Given: The battleship and enemy ships A
and B lie along a straight line. Neglect air
friction.
A battleship simultaneously fires two shells
at these two enemy ships.from a cliff 33.8 m high. At the level of the
sea, a rock sticks out a horizontal distance of
8.29 m.
The acceleration of gravity is 9.8 m/s2. With what minimum horizontal velocity
must the cliff divers leave the top of the cliff if
they are to miss the rock?
Answer in units of m/s.
1
Expert's answer
2019-10-10T09:44:16-0400

The time of flight:


"t=\sqrt{\frac{2h}{g}}"

The minimum horizontal velocity:


"v=\frac{d}{t}=d\sqrt{\frac{g}{2h}}"

"v=8.29\sqrt{\frac{9.8}{2(33.8)}}=3.16\frac{m}{s}"